### v1

@@ 3.4.6 Cellular Automata
Due to its importance in several areas of mathematics and physics, it is interesting to note that there has been an important series of papers devoted to understanding the exact difficulty (in the sense of “degree of uncomputability”) of several problems regarding spectra of infinite dimensional operators. One approach is classified via the so-called Solvability Complexity Index hierarchy introduced by Hansen (2011) and it is crucial to understand the possibility of using some undecidable problems as part of computer assisted proofs, as has been done for instance in the solution of Kepler’s conjecture Hales (2005); Hales et al. (2017). We will not delve into this topic here and just point to Colbrook and Hansen (2022) and the references therein, also for the long history of the problem of computing spectra of operators, which can be traced back at least to Schrödinger (1940).
@@ 5.1 Many-Body Systems
In this section, we deal with many-body spin systems, both classical and quantum. The mathematical framework to describe those systems is the following. We consider particles located on a $D$-dimensional lattice $\mathbb{Z}^{D}$. To the particle in position $i$ we associate a complex Hilbert space of dimension $d$ $\mathcal{H}_{i}=\mathbb{C}^{d}$, where we denote by $|1\rangle,\ldots|d\rangle$ its canonical (or computational) basis. $\Lambda_{L}$ will denote the cube $\mathbb{Z}^{D}\cap[-L,L]^{D}$. Examples of lattices $\Lambda_{L}$ for $L=2$ and $D=1$ and $D=2$ are:
@@
$H_{L}=\sum_{\langle i,j\rangle\in\Lambda_{L}}h_{i,j}+\sum_{i\in\Lambda_{L}}k_{i}$
where $\langle i,j\rangle$ denotes neighboring positions in the lattice, and $h_{i,j}$ (resp. $k_{i}$) is a hermitian operator acting on the Hilbert space $\mathcal{H}_{i}\otimes\mathcal{H}_{j}$ (resp. $\mathcal{H}_{i}$).
We will also restrict to the case of translational invariant interactions, with maybe the exception of the boundary terms $k_{i}$, where $i$ is an element of the boundary of $\Lambda_{L}$.
The Hamiltonian is called classical if the interactions $h_{i,j}$ and $k_{i}$ are diagonal in the computational basis.
The eigenvalues of $H_{L}$ are the energy levels of the system $\lambda_{0}(L)<\lambda_{1}(L)<\lambda_{2}(L)<\ldots$.
In many-body spin systems, one is often interested in $\lambda_{0}(L)$ and $\lambda_{1}(L)$. Their corresponding eigenvectors are called ground states and elementary excitations, respectively, and their difference $\gamma_{L}=\lambda_{1}(L)-\lambda_{0}(L)$ is called the spectral gap—typically considered in the limit $L\rightarrow\infty$.
@@
As we saw in Section 3.4.5, in a Wang tiling problem one is given a set of square-shaped tiles $\mathcal{T}$ with marked edges. Tiles cannot be rotated or reflected, just translated. For simplicity let us assume that the marks are colours. One is allowed to place one tile next to another one if the colours match at the common edge. With this simple rule, the tiling problem is whether one can tile, or not, the whole plane $\mathbb{Z}^{2}$, or a finite region of it like $\Lambda_{L}$. We draw here an example of a tile set with three elements and a valid tiling of size $2\times 5$:
It is simple to convert a tiling problem into a classical many-body system as follows. The number of tiles $|\mathcal{T}|$ becomes the dimension of the local Hilbert space, and the set of tiles corresponds precisely to the set of computational basis states $|t\rangle$. If on a pair of neighboring sites $i,j$, we define $P_{i,j}$ as the projector
@@
the classical Hamiltonian
@@
penalizes non-allowed configurations of tiles, leading to the following statement:
$\lambda_{0}(L)=0$ if there exists a valid tiling of $\Lambda_{L}$. In that case, the set of ground states corresponds to the set of all possible valid tilings. Otherwise $\lambda_{1}(L)\geq 1$.
Moreover, by adding boundary terms one can impose a particular configuration of tiles on the boundary of $\Lambda_{L}$.
Now, as specified for instance in the work of Robinson (1971), one can encode the evolution of any Turing Machine in a tiling problem so that, once the first row of the tiling is fixed, there is a unique way to complete the tiling, where the t-th row corresponds to the configuration of the Turing Machine at time t. By removing the tiles transitioning to the halting state, the problem of tiling becomes equivalent to the halting problem of a Turing Machine and hence undecidable.
Therefore, if we allow in this construction for an arbitrary Turing Machine and initialize it to the empty input, one obtains that it is undecidable to determine the ground state energy of a many-body classical system with translational invariant boundary conditions. The price to pay is that the local dimension (which in this case is the size of the Turing Machine) is unbounded.
One can remedy this by taking a Universal Turing Machine; but then one needs arbitrary (non-translationally invariant) boundary conditions that encode the input to the computation.
There are several ways in which one can improve this result that essentially dates back to the 80’s  [^21] The work of Kanter (1990) commented in Section 4 also has a similar flavour. in different directions (reducing the dimension, removing the boundary terms, going to 1D, focusing on particular relevant quantities of the ground and excited states,…). This is the content of the next sections.
@@ 5.1.1 Classical Systems
For classical systems, a central question is whether one can obtain undecidability for local dimension $2$, that is for the standard setup of $\{0,1\}$-valued classical spin systems, and whether the undecidability of the ground state energy translates to relevant quantities, like the magnetization. This is precisely the content of the contributions of Gu et al. (2009) and Gu and Perales (2012), where the evolution of a 1D cellular automata is encoded in the ground state of a 2D (non-translationally invariant) Ising model. Using such an encoding it is shown that the magnetization in the ground state, its degeneracy, and its correlation length are uncomputable quantities.
@@ 5.1.2 Quantum Systems
For quantum systems, it is possible to obtain undecidability in the nearest neighbor and pure translational invariant situation and with a fixed finite local dimension. In particular, for 2D quantum many-body systems of this kind, Cubitt et al. (2015b, 2022) show undecidability of the spectral gap, the ground state energy density and other physical quantities like the existence of power-law correlations in the ground state. Moreover, they show that undecidability of the spectral gap remains even if the considered Hamiltonians are arbitrary small perturbations of classical interactions and under the following promise: in the gapped case, the ground state is unique and product and the gap is of the size of the interaction strength; in the gapless case, the spectrum converges to $\mathbb{R}$ in the thermodynamic limit.
Building on the techniques developed by Cubitt et al. (2015b, 2022), this result was generalized later in several directions. In Bausch et al. (2020) it was shown to hold, with the same properties, in 1D systems. In Bausch et al. (2021) it was shown that, in 2D, phase diagrams are uncomputable, with both the gapped and gapless phases having non-zero measure but interleaved in an uncomputably complicated way. In Watson et al. (2022), it was shown that the renormalization group construction can also have an uncomputably complex behavior. Cubitt (2021) showed that there is a simple, direct argument that algorithmic undecidability implies the presence or absence of a spectral spectral gap is also axiomatically independent. Purcell et al. (2024) showed that even simple phase diagrams containing a single, first-order phase transition can be undecidable by constructing a system in which the phase transition occurs at an uncomputable number. [^22] In their construction, the phase transition occurs at Chaitin’s constant. This number, originally constructed by Chaitin (1975), can be thought of as the probability that the universal Turing Machine will halt on a randomly chosen input, where inputs are chosen with probability inverse-exponentially proportional to their lengths.
In a different direction, Shiraishi and Matsumoto (2021, 2020) have shown that the problem of thermalization in quantum many-body systems is also undecidable. In particular, even for the case of 1D nearest-neighbor interactions, they show that the long-time average of the expectation value of an observable (which can be taken as a spatial average of a single-site operator) is uncomputable.
@@ 5.1.3 Tensor Networks
Tensor Networks constitute a quantum information-inspired description of many-body quantum states since they aim to capture the entanglement pattern present in the state. They are one of the tools used to understand and simulate the low energy sector of quantum many-body Hamiltonians. We refer to Cirac et al. (2021) for a recent review.
Let us start with a directed graph $G=(V,E)$ and denote by $\deg^{-}(v)$ and $\deg^{+}(v)$ the ingoing and outgoing degrees of a vertex $v$ and by $\deg(v)=\deg^{-}(v)+\deg^{+}(v)$ the total degree of $v$. A tensor network (TN) over $G$ is defined by two ingredients. First, we associate to each edge $e$ of the graph a dimension $D_{e}$ and an index $i_{e}$ (those are named virtual indices). Second, for each vertex $v$ we select a complex-valued tensor $T^{v}$ with $\deg^{-}(v)$ covariant indices $i_{e}$, one for each incoming edge, and $\deg^{+}(v)+1$ contravariant indices $i_{e}$, one for each outcoming edge, plus an extra index $j_{v}\in\{1,\ldots,d\}$, called physical index. $D=\max_{e\in E}D_{e}$ is called the bond dimension of the tensor network. The tensor network $\mathcal{T}=(T^{v})_{v\in V}$ defines naturally a vector $\phi_{\mathcal{T}}\in\bigotimes_{v\in V}\mathbb{C}^{d_{v}}$, which is simply the result of contracting (i.e. summing over, in Einstein summation convention) all edge indices $i_{e}$ of the tensors $T_{v}$. Note that, if $N$ is the number of vertices in $G$, the number of parameters needed to describe the vector $\phi_{\mathcal{T}}\in\bigotimes_{v\in V}\mathbb{C}^{d_{v}}$ is bounded by
@@
which is polynomial in $N$ if the graph has bounded degree and $D={\rm poly}(N)$. This contrasts with the exponential number $\prod_{v\in V}d_{v}$ required to describe an arbitrary vector $\phi\in\bigotimes_{v\in V}\mathbb{C}^{d_{v}}$. Therefore tensor networks can be understood as efficient ways of describing (families of) vectors in large tensor product spaces. By a sequence of major results of Hastings (2006, 2007), improved later by several authors (see the review by Cirac et al. (2021)) this compressed description is still accurate enough to accurately approximate thermal states of locally interacting quantum systems, and the ground and excited states of 1d locally interacting quantum systems. This is why variational algorithms that optimize within the manifold of Tensor Network states have proven to be highly accurate in these cases.
@@
The most basic example of a TN is a Matrix Product State (MPS), also known in the numerical analysis community as a Tensor Train (Cichocki et al. (2016, 2017)), which has a line as its associated graph. The graphical representation of a translationally invariant MPS, where the tensor is the same in every vertex, is then
@@
With this definition, it is clear that the contraction of the translationally invariant TN defined by $T$ is $0$ in a given region if and only if there is a valid tiling of that region, which makes the problem of tensor contraction of the infinite translationally invariant TN (distinguishing whether it is 00 or $\geq 1$) undecidable.
Variants of this idea, with other undecidable problems as primitive instead of tiling, were reported by Morton and Biamonte (2012), where the first undecidable problems in the context of TN were analyzed.
This mapping from tiling problems to TN was explicitly used by Scarpa et al. (2020). They show that many crucial properties for the use of PEPS in quantum many-body problems, both analytically and numerically, are indeed undecidable in the thermodynamic limit. The most paradigmatic one is whether a given PEPS is invariant under a global on-site symmetry. That is, given a group $G$ and an associated representation $U_{v}(g)$ for each vertex $v$ of the graph $G$, decide whether $\bigotimes_{v}U_{v}(g)\phi_{\mathcal{T}}=\phi_{\mathcal{T}}$ for all $g\in G$.
@@
This problem has been known to be hard in practice for a long time (Verstraete et al. (2004)) and the work-around solution was to take arbitrary Matrix Product Operators $\Theta$ and use $\Theta\Theta^{\dagger}$ as the variational family in the optimization algorithms. Note that, $\Theta\Theta^{\dagger}$ is also a Matrix Product Operator as one can easily see from the graphical notation.
@@
In Kliesch et al. (2014) it was finally proven that deciding whether a given Matrix Product Operator is positive or not for all system sizes is undecidable. Given the formula (2), it does not come as a surprise that the result was obtained by a reduction to matrix mortality problems (see section 3.4.3).
The result of Kliesch et al. (2014) was used later by De las Cuevas et al. (2016) to prove the existence of Matrix Product Density Operators without a local purification. The proof uses the quantum version of Wielandt’s inequality of Sanz et al. (2010) to show that, in case a local purification exists, there is an algorithm that finds it. Therefore, if all Matrix Product Density Operators had a local purification, checking whether a given Matrix Product Operator was positive semidefinite would be decidable, contradicting the result in Kliesch et al. (2014).

### v2

@@ 3.4.6 Cellular Automata
Due to its importance in several areas of mathematics and physics, it is interesting to note that there has been an important series of papers devoted to understanding the exact difficulty (in the sense of “degree of uncomputability”) of several problems regarding spectra of infinite dimensional operators. One approach is classified via the so-called Solvability Complexity Index hierarchy introduced by Hansen (2011) [121] and it is crucial to understand the possibility of using some undecidable problems as part of computer assisted proofs, as has been done for instance in the solution of Kepler’s conjecture [122, 123]. We will not delve into this topic here and just point to the work by Colbrook and Hansen (2022) [124] and the references therein, also for the long history of the problem of computing spectra of operators, which can be traced back at least to Schrödinger (1940) [125].
@@ 5.1 Many-Body Systems
In this section, we deal with many-body spin systems, both classical and quantum. The mathematical framework to describe those systems is the following. We consider particles located on a $D$-dimensional lattice $\mathbb{Z}^{D}$. To the particle in position $i$ we associate a $d$-dimensional complex Hilbert space $\mathcal{H}_{i}=\mathbb{C}^{d}$, where we denote by $|1\rangle,\ldots|d\rangle$ its canonical (or computational) basis. $\Lambda_{L}$ will denote the cube $\mathbb{Z}^{D}\cap[-L,L]^{D}$. Examples of lattices $\Lambda_{L}$ for $L=2$ and $D=1$ and $D=2$ are:
@@
$H_{L}=\sum_{\langle i,j\rangle}h_{i,j}+\sum_{i\in\Lambda_{L}}k_{i}$ (2)
where $\langle i,j\rangle$ denotes neighboring positions in the lattice $\Lambda_{L}$, and $h_{i,j}$ (resp. $k_{i}$) is a Hermitian operator acting on the Hilbert space $\mathcal{H}_{i}\otimes\mathcal{H}_{j}$ (resp. $\mathcal{H}_{i}$).
We will also restrict to the case of translational invariant interactions, with maybe the exception of those single site terms $k_{i}$ for which $i$ is an element of the boundary of $\Lambda_{L}$. That is, in the bulk of $\Lambda_{L}$, all interaction terms $k_{i}$ will be just translations of a fixed Hermitian operator $k$ acting on $\mathbb{C}^{d}$, whereas for sites $i$ on the boundary, the corresponding $k_{i}$ will sometimes be allowed to vary from site to site. For the two-body terms the operators $h_{i,j}$ and $h_{i^{\prime},j^{\prime}}$ coincide if $i-i^{\prime}=j-j^{\prime}$. For instance, in 2D all horizontal interaction terms are translations of a fixed Hermitian operator $h^{\rm row}$ acting on $\mathbb{C}^{d}\otimes\mathbb{C}^{d}$ and all vertical interaction terms are translations of a fixed hermitian operator $h^{\rm col}$ acting on $\mathbb{C}^{d}\otimes\mathbb{C}^{d}$.
The Hamiltonian is called classical if the interactions $h_{i,j}$ and $k_{i}$ are diagonal in the computational basis. The eigenvalues of $H_{L}$ are the energy levels of the system $\lambda_{0}(L)<\lambda_{1}(L)<\lambda_{2}(L)<\ldots$. In many-body spin systems, one is often interested in $\lambda_{0}(L)$ and $\lambda_{1}(L)$. Their corresponding eigenvectors are called ground states and elementary excitations, respectively, and their difference $\gamma_{L}=\lambda_{1}(L)-\lambda_{0}(L)$ is called the spectral gap—typically considered in the limit $L\rightarrow\infty$.
@@
As we saw in Section 3.4.5, in a Wang tiling problem one is given a set of square-shaped tiles $\mathcal{T}$ with marked edges. Tiles cannot be rotated or reflected, just translated. For simplicity let us assume that the marks are colours. One is allowed to place one tile next to another one if the colours match at the common edge. With this simple rule, the tiling problem is whether one can tile, or not, the whole plane $\mathbb{Z}^{2}$ or, equivalently, due to Wang’s extension theorem [84], the set of finite regions $\Lambda_{L}$. One can consider other versions of the tiling problem, by restricting to infinite subsets of $\mathbb{Z}^{2}$, like the upper-half plane. We draw here an example of a tile set with three elements and a valid tiling of size $2\times 5$:
It is simple to convert a tiling problem into a classical many-body system as follows. The number of tiles $|\mathcal{T}|$ becomes the dimension of the local Hilbert space, and the set of tiles corresponds precisely to the set of computational basis states $|t\rangle$. On a pair of neighboring sites $i,j$, we say that a pair of tiles $(t_{i},t_{j})\in\mathcal{T}\times\mathcal{T}$ is not allowed if they have different colors in their common edge, that is, if they cannot be placed together in a valid tiling. This notation allows us to define, on each pair of neighboring sites $i,j$, the projector
@@
which gives rise to the (translational invariant) classical Hamiltonian
@@
that penalizes non-allowed configurations of tiles. This construction leads to the following statement for the ground state energy $\lambda_{0}(L)$ of $H_{L}$:
$\lambda_{0}(L)=0$ if there exists a valid tiling of $\Lambda_{L}$. In that case, the set of ground states corresponds to the set of all possible valid tilings. Otherwise $\lambda_{0}(L)\geq 1$.
Since, as we saw in Section 3.4.5, the tiling problem is undecidable, the same holds true for the problem of determining whether the ground state energy $\lambda_{0}(L)=0$ for all $L$.
Note that in this reduction from the tiling problem to the ground state energy problem, the local dimension (which is the size of the tiling set) is unbounded. However, one can easily obtain the result for a fixed local dimension at the price of adding boundary terms which break translational invariance. For that we consider the problem of tiling the upper-half plane where one fixes the tiles in the first row. There is a straightforward way to encode the evolution of any Turing Machine in a tiling problem so that, once the first row of the tiling is fixed, there is a unique way to complete the tiling, wherein the t-th row corresponds to the configuration of the Turing Machine at time $t$ (see e.g. the discussion in Robinson (1971) [80]). By removing the tiles representing transitioning to the halting state, the problem of tiling becomes equivalent to the halting problem of a Turing Machine and hence undecidable. Therefore, if we consider in this construction a Universal Turing Machine, initialize it to any input and consider the corresponding tiling to Hamiltonian reduction, it is not difficult to conclude that it is undecidable to determine whether $\lambda_{0}(L)=0$ for a sequence of many-body classical Hamiltonians $H_{L}$ as in equation (2), which are translational invariant in the bulk, have a fixed local dimension, but have non-translational invariant boundary conditions.
There are several ways in which one can improve this result that essentially dates back to the 80’s  [^22] The work [81] mentioned in Section 4 also has a similar flavour. in different directions (reducing the dimension, removing the boundary terms, going to 1D, focusing on particular relevant quantities of the ground and excited states,…). This is the content of the next sections.
@@ 5.1.1 Classical Systems
For classical systems, a central question is whether one can obtain undecidability for local dimension $2$, that is for the standard setup of $\{0,1\}$-valued classical spin systems, and whether the undecidability of the ground state energy translates to relevant quantities, like the magnetization. This is precisely the content of the contributions of Gu et al. (2009) [96] and Gu and Perales (2012) [97], where the evolution of a 1D cellular automata is encoded in the ground state of a 2D (non-translationally invariant) Ising model. Using such an encoding it is shown that the magnetization in the ground state, its degeneracy, and its correlation length are uncomputable quantities.
@@ 5.1.2 Quantum Systems
For quantum systems, it is possible to obtain undecidability in the nearest neighbor and pure translational invariant situation and with a fixed finite local dimension. In particular, for 2D quantum many-body systems of this kind, Cubitt et al. (2015, 2022) [82, 120] showed undecidability of the spectral gap, the ground state energy density and other physical quantities like the existence of power-law correlations in the ground state. Moreover, they showed that undecidability of the spectral gap remains even if the considered Hamiltonians are arbitrary small perturbations of classical interactions and under the following promise: in the gapped case, the ground state is unique and a product state and the gap is of the size of the interaction strength; in the gapless case, the spectrum converges to $\mathbb{R}$ in the thermodynamic limit.
Building on the techniques developed in [82, 120], this result was generalized later in several directions. It was shown to hold, with the same properties, in 1D systems by Bausch et al. (2020) [126]. Bausch et al. (2021) [127] showed that, in 2D, phase diagrams are uncomputable, with both the gapped and gapless phases having non-zero measure but interleaved in an uncomputably complicated way. Watson et al. (2022) [128] showed that the renormalization group construction can also have an uncomputably complex behavior. Cubitt (2021) [129] showed that there is a simple, direct argument that algorithmic undecidability implies the presence or absence of a spectral gap is also axiomatically independent. Purcell et al. (2024) [45] showed that even simple phase diagrams containing a single, first-order phase transition can be undecidable by constructing a system in which the phase transition occurs at an uncomputable number. [^23] In their construction, the phase transition occurs at Chaitin’s constant. This number, originally constructed by Chaitin [130], can be thought of as the probability that the universal Turing Machine will halt on a randomly chosen input, where inputs are chosen with probability inverse-exponentially proportional to their lengths.
In a different direction, Shiraishi and Matsumoto (2021, 2020) [131, 132] have shown that the problem of thermalization in quantum many-body systems is also undecidable. In particular, even for the case of 1D nearest-neighbor interactions, they show that the long-time average of the expectation value of an observable (which can be taken as a spatial average of a single-site operator) is uncomputable.
@@ 5.1.3 Tensor Networks
Tensor Networks constitute a quantum information-inspired description of many-body quantum states since they aim to capture the entanglement pattern present in the state. They are one of the tools used to understand and simulate the low energy sector of quantum many-body Hamiltonians. We refer to [133] for a recent review.
Let us start with a directed graph $G=(V,E)$ and denote by $\deg^{-}(v)$ and $\deg^{+}(v)$ the ingoing and outgoing degrees of a vertex $v$ and by $\deg(v)=\deg^{-}(v)+\deg^{+}(v)$ the total degree of $v$. A tensor network (TN) over $G$ is defined by two ingredients. First, we associate to each edge $e$ of the graph a dimension $D_{e}$ and an index $i_{e}$ (called a virtual index). Second, for each vertex $v$ we select a complex-valued tensor $T^{v}$ with $\deg^{-}(v)$ covariant indices $i_{e}$, one for each incoming edge, and $\deg^{+}(v)+1$ contravariant indices $i_{e}$, one for each outcoming edge, plus an extra index $j_{v}\in\{1,\ldots,d\}$, called the physical index. The dimension $D=\max_{e\in E}D_{e}$ is called the bond dimension of the tensor network. The tensor network $\mathcal{T}=(T^{v})_{v\in V}$ defines naturally a vector $\phi_{\mathcal{T}}\in\bigotimes_{v\in V}\mathbb{C}^{d_{v}}$, which is simply the result of contracting (i.e. summing over, in Einstein summation convention) all edge indices $i_{e}$ of the tensors $T_{v}$. Note that, if $N$ is the number of vertices in $G$, the number of parameters needed to describe the vector $\phi_{\mathcal{T}}\in\bigotimes_{v\in V}\mathbb{C}^{d_{v}}$ is bounded by
@@
which is polynomial in $N$ if the graph has bounded degree and $D={\rm poly}(N)$. This contrasts with the exponential number $\prod_{v\in V}d_{v}$ required to describe an arbitrary vector $\phi\in\bigotimes_{v\in V}\mathbb{C}^{d_{v}}$. Therefore tensor networks can be understood as efficient ways of describing (families of) vectors in large tensor product spaces. By a sequence of major results of Hastings [134, 135], improved later by several authors (see the review [133]) this compressed description is still accurate enough to accurately approximate thermal states of locally interacting quantum systems, and the ground and excited states of 1d locally interacting quantum systems. This is why variational algorithms that optimize within the manifold of Tensor Network states have proven to be highly accurate in these cases.
@@
The most basic example of a TN is a Matrix Product State (MPS), also known in the numerical analysis community as a Tensor Train [136, 137], which has a line as its associated graph. The graphical representation of a translationally invariant MPS, where the tensor is the same in every vertex, is then
@@
With this definition, it is clear that the contraction of the translationally invariant TN defined by $T$ is not equal to zero in a given region if and only if there is a valid tiling of that region, which makes the problem of tensor contraction of the infinite translationally invariant TN (distinguishing whether it is 00 or $\geq 1$) undecidable.
Variants of this idea, with Word Problem as primitive instead of tiling, were reported by Morton and Biamonte (2012) [73], where the first undecidable problems in the context of TN were analyzed.
This mapping from tiling problems to TN was explicitly used by Scarpa et al. (2020) [83]. They show that many crucial properties for the use of PEPS in quantum many-body problems, both analytically and numerically, are indeed undecidable in the thermodynamic limit. The most paradigmatic one is whether a given PEPS is invariant under a global on-site symmetry. That is, given a group $G$ and an associated representation $U_{v}(g)$ for each vertex $v$ of the graph $G$, decide whether $\bigotimes_{v}U_{v}(g)\phi_{\mathcal{T}}=\phi_{\mathcal{T}}$ for all $g\in G$.
@@
This problem has been known to be hard in practice for a long time [138] and the work-around solution was to take arbitrary Matrix Product Operators $\Theta$ and use $\Theta\Theta^{\dagger}$ as the variational family in the optimization algorithms. Note that, $\Theta\Theta^{\dagger}$ is also a Matrix Product Operator as one can easily see from the graphical notation.
@@
Kliesch et al. (2014) [68] finally proved that deciding whether a given Matrix Product Operator is positive or not for all system sizes is undecidable. Given the formula (3), it does not come as a surprise that the result was obtained by a reduction to matrix mortality problems (see section 3.4.3).
This result was used later by De las Cuevas et al. (2016) [69] to prove the existence of Matrix Product Density Operators without a local purification. The proof uses the quantum version of Wielandt’s inequality [139] to show that, in case a local purification exists, there is an algorithm that finds it. Therefore, if all Matrix Product Density Operators had a local purification, checking whether a given Matrix Product Operator was positive semidefinite would be decidable, contradicting the result in [68].